Hindi Builders Inc. produces three products: A, B, and C. The following information is presented for the three products:

Fixed Cost

$       687,400  

Product A

Product B

Product C

Units produced

80  

120  

200  

Price Per Unit

$               320  

$         1,800  

$         2,800  

Variable Cost Per Unit

$               160  

$            800  

$         1,500  

Required:

  1. Calculate the contribution margin for each product

  1. Calculate the break-even point in units of the three products A, B, and C combination based on the sales mix percentage

  1. Please give suggestions to the decision makers about how to increase profit based on the CVP analysis.
Answer #1

Answers

  • All working forms part of the answer
  • Requirement 1: Contribution margin calculation

Product A

Product B

Product C

A

Price per unit

$320

$1,800

$2,800

B

Variable cost per unit

$160

$800

$1,500

C = A-B

Contribution margin per unit = Answer

$160

$1,000

$1,300

  • Requirement 2

A

B

C

Total

A

Unit produced [Sales Mix]

80

120

200

400

B = A/400 units

Sales Mix % of total sales

20%

30%

50%

100%

C

Contribution margin per unit = Answer

$160

$1,000

$1,300

D = B x C

Weighted Average CM per unit

$32

$300

$650

$982

E

Total Fixed Cost

$687,400

F = E/D

Total Break even units

                         700

G = F x B

Break even of products = Answers

                                    140

                           210

                     350

  • Requirement 3

Suggestions:
>As it can be seen that Actual units produced (400 units) are far LESS than break even units required (700 units). Hence, first objective would be to achieve break even and minimise the loss.
>Product C’s Breakeven is 350 units, BUT only 200 units have been produced. These should be produced and sold atleast more than break even point of 350 units.
>Product B’s breakeven is 210 units BUT only 120 units have been produced. This should be produced in quantity of 210 units or more.
>Product A’s Weighted average CM per unit is the lowest and its breakeven targets are not achieved of 140 units.
>In order to increase profits, the units are to be produced and sold over and above the quantities of Break even.